OpenStudy (anonymous):

find the x intercept and y intercept of the line. -9x+8y=14

OpenStudy (nikato):

to find the x interceot, make y=0 and solve for x for y intercept, make x=0 and solve for y

OpenStudy (anonymous):

im still having trouble

OpenStudy (nikato):

ok, so lets find the x-intercept first to do this we plug in 0 for y int he equation so we have -9x + 8(0)= 14 can you solve for x?

OpenStudy (bookworm14):

if you substitute x with zero you will get the new equation -9(0)+8y=14 then simplify 8y=14 in order to do this divide 14/8 resulting in 1 and 3/4 or 14/8 now substitute zero for y resulting in the new equation -9x+8(0)=14 now simplify -9x=14 in order to do this you divide 14/ -9 resulting in -1 and 5/9 or -1.555555556 Hope This Helped !!